Please solve it by hand or typing, don't use EXCEL.

1. XYZ stock price and dividends history are as follows Year Beginning-of-year price Dividend paid at year- $100 $110 $90 $95 end $4 $4 $4 $4 2005 2006 2007 2008 An investor buys three shares of XYZ at the beginning of 2005, buys another two shares at the beginning of 2006, sells one share at the beginning of 2007, and sells all four remaining shares at the beginning of 2008. Hint: When the investor buys three shares at the beginning of 2005, she cannot receive dividends paid at the end of 2004. She can receive dividend payments on these three shares at the end of 2005 A. Computer HPR for each period (05-06, 06-07, and 07-08) (6pts) B. Compute the arithmetic return. (3pts) C. Compute the geometric return. (3pts) D. Computer the dollar-weighted rate of return. Draw the time line to show cash flows. Write down the formula and calculate the dollar-weighted rate of return. The time line includes four dates corresponding to the turns of the year for January 1, 2005 to January 1, 2008. (8pts)
Answer #1

YEAR SHARE(BUY/ SALE) PRICE DIVIDEND

2005 3 $100    $4

2006 2 $110    $4

2007 (1) $90 $4

2008 (4) $95 $4

0

NOTE: WE ASSUMED THAT DIVIDEND AT THE END OF THE YEAR IS PER SHARE $ 4

1. HOLDING PERIODS RETURN

YEAR   

2005-2006 3 * $4 = $12

2006-2007 5 * $4 = $20

2007-2008 4 *$4 = $16

2007-2008 $90- $105 = (15) SO HOLDING PERIODS RETURN FOR 2007-08 $16- $15 = $1

2008-2009 ($95- $105)*3 = (30) SO HOLDING PERIODS RETURN FOR 2008-09 $15

TOTAL HPR = $20+$16+$1-$30 = $7

W.K. AVERAGE PRICE FOR SELLING AT THE BEGINNING OF THE YEAR 2007-2008

100+110 / 2 = 105

2. ARITHMETIC RETURN:-

BUYING AVG. COST = 105 * 5 = 525

SELLING = 90+95/2 * 5   = 462.5

NET = 62.5

+ DIVIDEND RECIVED = $48

NET PROFIT = (14.5)